# On Mean Field Games in Infinite Dimension

### v1

@@ 2.3. The operator $A$ and the generated semigroup
    It can be verified from (2.7) that Assumption 2.1 implies typical conditions on the operator $A$ ensuring existence and uniqueness of mild solutions of equations in infinite dimensional spaces (see [19, Ch. 4]), that is:
@@ 2.4. The probabilistic framework and the reference SDE
Under Assumption 2.1, the stochastic convolution (see Remark 2.2(ii) and [18, Ch. 5])
@@
called the Ornstein-Uhlenbeck semigroup. We will use it later to define the concept of a mild solution to (HJB).
@@
When $w$ does not depend on $t$, the result is proved [14, Th. 13] and the proof works in the same way if $w$ depends on $t$. We just need to explain how our setup fits into the framework of [14].
Given our filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathsf{P})$ with the Wiener process $W$, let $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ be another complete atomless probability space. For any any $m_{0}\in\mathcal{P}_{1}(H)$, since $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ is nonatomic, there exists an $\mathcal{F}_{1}$-measurable random variable $\xi_{0}:\Omega_{1}\rightarrow X$ such that $m_{0}=\mathcal{L}(\xi_{0})$ (see, e.g., [9, Lemma 5.29]). We define a new probability space $(\Omega_{2},\mathcal{F}_{2},\mathsf{P}_{2})$ as follows:
@@
Thus, the law of $X_{0}$ in the new probability space is the same. This puts us in the framework of [14] and we can now proceed as in the proof of [14, Th. 13].
@@
    Using the definition of mild solution, the fact that $|e^{tA}|\leq 1$ and the estimate on the stochastic convolution $W_{t}^{A}$ [19, Prop. 1.144], we have, for some $c_{1}>0$,
@@
      The map $[0,T]\rightarrow L^{2}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous (see [18, Th. 5.2]); hence, also the map $[0,T]\rightarrow L^{1}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous. Hence, since $[0,T]$ is compact, there exists a modulus of continuity $\hat{\omega}_{A}$ such that, for every $0\leq s\leq t\leq T$,
@@ 3. The Fokker-Planck Equation: Weak Solutions
In this section, we study (FP) when the term $\mathcal{H}_{p}(x,Dv(t,x),m(t))$ is frozen. For that, we fix $w\in C_{b}((0,T)\times H;H$) and consider the linear Fokker-Planck equation in the space $\mathcal{P}_{1}(H)$
@@
We will interpret (3.1) in a weak sense. For that, we define the set of test functions
(3.2) $\mathcal{D}_{T}=\Big{\{}\varphi\in C_{b}^{1,2}([0,T]\times H):\ A^{*}D\varphi(t,x)\in C_{b}([0,T]\times H;H),\ \ D^{2}\varphi\in C_{b}([0,T]\times H;\mathcal{L}_{1}(H))\Big{\}}.$
To provide a rigorous definition of solution to (3.1), we introduce the following operator [^1] Note that, as explained in [19, Section B.5.2], the operator $L_{0}$ is closable in a suitable weak sense (so called $\mathcal{K}$-closable) and its closure (called $\mathcal{K}$-closure there) is the generator (again in a suitable weak sense) of the semigroup $R_{t}$ in (2.10).
(3.3) $(L_{0}\phi)(x)=\langle x,A^{*}D\phi(x)\rangle+\frac{1}{2}\mbox{Tr}\,[D^{2}\phi(x)],$
defined on
(3.4) $D(L_{0})$ =$=\Big{\{}\phi\in UC^{2}_{b}(H):\ D^{2}\phi\in UC_{b}(H;\mathcal{L}_{1}(H)),\ A^{*}D\phi\in UC_{b}(H;H)\Big{\}}.$
The formal operators $L^{*}$ and ${\mathbf{div}}$ in (3.1) are interpreted similarly as for a Fokker-Planck equation in the case $H=\mathbb{R}^{n}$ when using the class of test functions $\mathcal{D}_{T}$. The definition of weak solution to (3.1) given below is the same as the one of [5]; however, we use a larger class of test functions.
@@
A map $m:[0,T]\rightarrow\mathcal{P}_{1}(H)$ such that $t\mapsto m(t)(I)$ is measurable for each $I\in\mathcal{B}(H)$ [^2] That is, in the terminology of [5], $m$ is a probability kernel. is called a weak solution to (3.1) if $m(0)=m_{0}$ and
@@
for every $t\in[0,T]$ and $\varphi\in\mathcal{D}_{T}$.
@@
This is just an application of Dynkin’s formula for regular functions: see, e.g, in [19, Prop. 1.164], by interpreting $(w(s,X^{w}_{s}))_{s\in[0,T]}$ as the process $(f_{s})_{s\in[0,T]}$ there.
@@
The claim follows from [5, Th. 4.1] once we prove that a solution in our sense is also a solution in the sense of [5, Th. 4.1]. To do that we first notice that, in [5, Formula (1.5)], the definition of solution is given using the set test functions (called $\mathcal{E}_{A}(H)$ there) which is the linear span of the real parts of the functions of the form
@@
Given the above, we observe that such a set is clearly contained in the set $\mathcal{D}_{T}$ defined in (3.2). Finally, we point out that in [5] the Kolmogorov operator $L$ is defined differently (corresponding to our $\partial_{t}+L_{0}-\langle w,D\rangle$); still, equality (1.5) in [5] corresponds to our (3.5).
@@
By Proposition 2.5(ii), our solution provided by Proposition 3.2 belongs to the class $\mathcal{S}_{4}$ and the uniqueness follows from Proposition 3.3.
@@ 4. The HJB Equation: Mild Solutions
In this section we deal with equation (HJB) for a given fixed $m\in\mathcal{S}$. We introduce the concept of a mild solution to (HJB) (see [19, Section 4.4.1.2]).
@@
    This part is contained in [19, Theorem 4.90]. [^3] More precisely, it is enough to take in that theorem, $m=0$, $\gamma_{G}(t)=t^{-\gamma}$, $G=\mathbf{Id}$. We warn that $G$ in this footnote refers to the operator used in [19, Theorem 4.90], which has nothing to do with our function $G$.
@@
    Using the smoothing properties of the Ornstein-Uhlenbeck semigroup (see e.g. [19, Theorem 4.37]), we have the representation of the gradients
@@
    Hence, fixing $t\in[0,T)$ and considering [19, Th. 4.37] and Assumption 2.6, we have, for some $C>0$,
@@
    By a suitable generalization of Gronwall’s Lemma (see [19, Prop. D.30]), we get
@@ 5. The MFG System: Existence of Solutions
    $\sup_{t\in[0,T],\,x\in H}\int_{0}^{t}\langle e^{(t-s)A}w(s,x),\mathbf{e}_{n}\rangle^{2}\mathrm{d}s\leq\alpha_{n}\cdot\sup_{s\in(0,T),\,x\in H}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\ \leq\ \alpha_{n}{\|w\|^{2}_{\infty}},\ \ \ \forall n\in\mathbb{N}.$
@@
    $\int_{H}\sup_{t\in[0,T]}\langle e^{tA}x,\mathbf{e}_{i}\rangle^{2}m_{0}(\mathrm{d}x)=\int_{H}\sup_{t\in[0,T]}\langle x,e^{tA}\mathbf{e}_{n}\rangle^{2}m_{0}(\mathrm{d}x)$
@@
    $=\sup_{t\in[0,T],\,x\in H}\ \int_{0}^{t}e^{2(t-s)\lambda_{n}}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\mathrm{d}s\leq\sup_{s\in(0,T),\,x\in H}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\cdot\sup_{t\in[0,T]}\int_{0}^{t}e^{2(t-s)\lambda_{n}}\mathrm{d}s$
@@
Observe that, by Lemma A.1 and by [38, Remark 7.13 (ii)], the metric $\mathbf{d}_{1}$ metrizes on $\mathcal{R}_{m_{0}}$ the topology of weak convergence in $\mathcal{P}(H)$. In view of Proposition 5.2, we also define
@@
We will use the fact (see, e.g., [38, Rem. 7.5(i)] or [9, Cor. 5.4]) that for $\mu,\nu\in\mathcal{P}_{1}(H)$
@@
6. The MFG System: Existence and Uniqueness of Solutions
In this section, we establish the uniqueness of solutions for our MFG system (HJB)–(FP). It is well known, even in finite-dimensional state spaces (see, e.g., [8, Rem. 1.4, p.32]), that uniqueness of a solution for this system cannot generally be guaranteed. A typical approach to ensure uniqueness is to impose certain monotonicity conditions on the data. One of the most common monotonicity conditions is the Lasry-Lions monotonicity condition (see, for example, [30, 31, 32] and [8, Section 1.3.2]). For other monotonicity conditions that ensure uniqueness in finite-dimensional MFG systems, readers can refer to [28, 34] and the references therein. Another typical feature of uniqueness results is the separability of the Hamiltonian function $\mathcal{H}$ (as assumed in Assumption 6.1(i) here). Well-posedness of MFG systems without separability has been studied in [34]. Here, we assume separability and use a variant of the Lasry-Lions monotonicity condition from [8, Theorem 1.4]. Our assumptions are as follows.
@@ 6. The MFG System: Existence and Uniqueness of Solutions
      $\int_{H}[F(x,\mu_{1})-F(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)>0\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{1}(H),\;\mu_{1}\neq\mu_{2};$
@@
      $\int_{H}[G(x,\mu_{1})-G(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)=0\quad\Longrightarrow\quad G(\cdot,\mu_{1})=G(\cdot,\mu_{2});$
@@
Uniqueness. The proof is an adaptation to our Hilbert space setting of a typical proof of uniqueness in finite dimensional spaces (see, e.g., [8, Theorem 1.4]).
@@
$\bar{v}(t,x)$ $=R_{T-t}[G(\cdot,m_{1}(T))-G(\cdot,m_{2}(T))](x)-\int_{t}^{T}R_{s-t}[\mathcal{H}^{0}(\cdot,Du_{1}(s,\cdot))-\mathcal{H}^{0}(\cdot,Du_{2}(s,\cdot))](x)\mathrm{d}s$
$\ \ \ -\int_{t}^{T}R_{s-t}[F(\cdot,m_{1}(s))-F(\cdot,m_{2}(s)](x)\mathrm{d}s,$
@@
$-\int_{0}^{T}\left(\int_{H}\left[\langle\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x)),\,D\varphi(s,x)\rangle\right]m_{1}(s,\mathrm{d}x)-\int_{H}\left[\langle\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x)),\,D\varphi(s,x)\rangle\right]m_{2}(s,\mathrm{d}x)\right)\mathrm{d}s.$
@@
by the Dominated Convergence Theorem. The convergence of the other terms follows directly from the Dominated Convergence Theorem, since the functions involved are uniformly bounded. Therefore, passing to the limit in (6.3), we get
@@
    If case (a) of Assumption 6.1(iv) holds, we have $m_{1}=m_{2}$ and then uniqueness of mild solutions of equation (HJB) gives $v_{1}=v_{2}$.
@@
    If case (b) of Assumption 6.1(iv) holds, we first use uniqueness of mild solutions of equation (HJB) which holds since $F(x,m_{1}(s))=F(x,m_{2}(s))$ and $G(x,m_{1}(T))=F(x,m_{2}(T))$ to obtain $v_{1}=v_{2}$ and then conclude that $m_{1}=m_{2}$ by uniqueness of weak solutions of (3.5) with $w(s,x):=\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x))$.
@@
    If case (c) of Assumption 6.1(iv) holds, we first obtain $\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x))$, $m_{1}(s)$ and $m_{2}(s)$ a.e. for $s\in[0,T]$. This implies $m_{1}=m_{2}$ by uniqueness of weak solutions of (3.5) with $w(s,x)=\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))$ (or equivalently with $w(s,x)=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x))$) and then conclude that $v_{1}=v_{2}$ using uniqueness of mild solutions of equation (HJB).

### v2

@@ 2.3. The operator $A$ and the generated semigroup
    It can be verified from (2.7) that Assumption 2.1 implies typical conditions on the operator $A$ ensuring existence and uniqueness of mild solutions of equations in infinite dimensional spaces (see [21, Ch. 4]), that is:
@@ 2.4. The probabilistic framework and the reference SDE
Under Assumption 2.1, the stochastic convolution (see Remark 2.2(ii) and [20, Ch. 5])
@@
called the Ornstein-Uhlenbeck semigroup. Above, $\mathcal{N}(a,Q)$ is the Gaussian measure in $H$ with mean $a\in H$ and covariance operator $Q\in L^{1}_{+}(H)$ (see, e.g., [19, Ch.2, Sec.3]). We will use it later to define the concept of a mild solution to (HJB).
@@
When $w$ does not depend on $t$, the result is proved [16, Th. 13] and the proof works in the same way if $w$ depends on $t$. We just need to explain how our setup fits into the framework of [16].
Given our filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathsf{P})$ with the Wiener process $W$, let $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ be another complete atomless probability space. For any any $m_{0}\in\mathcal{P}_{1}(H)$, since $(\Omega_{1},\mathcal{F}_{1},\mathsf{P}_{1})$ is nonatomic, there exists an $\mathcal{F}_{1}$-measurable random variable $\xi_{0}:\Omega_{1}\rightarrow X$ such that $m_{0}=\mathcal{L}(\xi_{0})$ (see, e.g., [11, Lemma 5.29]). We define a new probability space $(\Omega_{2},\mathcal{F}_{2},\mathsf{P}_{2})$ as follows:
@@
Thus, the law of $X_{0}$ in the new probability space is the same. This puts us in the framework of [16] and we can now proceed as in the proof of [16, Th. 13].
@@
    Using the definition of mild solution, the fact that $|e^{tA}|\leq 1$ and the estimate on the stochastic convolution $W_{t}^{A}$ [21, Prop. 1.144], we have, for some $c_{1}>0$,
@@
      The map $[0,T]\rightarrow L^{2}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous (see [20, Th. 5.2]); hence, also the map $[0,T]\rightarrow L^{1}(\Omega,\mathsf{P})$, $t\mapsto W^{A}_{t}$ is continuous. Hence, since $[0,T]$ is compact, there exists a modulus of continuity $\hat{\omega}_{A}$ such that, for every $0\leq s\leq t\leq T$,
@@ 3. The Fokker-Planck Equation: Weak Solutions
In this section, we study (FP) when the term $\mathcal{H}_{p}(x,Dv(t,x),m(t))$ is frozen. So, given $w\in C_{b}((0,T)\times H;H$), we consider the linear Fokker-Planck equation in the space $\mathcal{P}_{1}(H)$
@@
We will interpret (3.1) in a weak sense, following the approach of [6]. For that, we define the set of test functions
[…]
(3.2) $\ D^{2}\varphi\in C_{b}([0,T]\times H;\mathcal{L}_{1}(H))\Big{\}}.$
To provide a rigorous definition of solution to (3.1), we introduce the operator
(3.3) $(L_{0}\phi)(x)=\langle x,A^{*}D\phi(x)\rangle+\frac{1}{2}\mbox{Tr}\,[D^{2}\phi(x)].$
We notice that $L_{0}\varphi(t,x):=[L_{0}\varphi(t,\cdot)](x)$ is well defined for every $\varphi\in\mathcal{D}_{T}$ and $t\in[0,T]$. The formal operators $L^{*}$ and ${\mathbf{div}}$ in (3.1) are interpreted similarly as for the Fokker-Planck equation in the case $H=\mathbb{R}^{n}$ when we apply test functions in $\mathcal{D}_{T}$. The definition of a weak solution to (3.1) given below is similar to the one of [6]; however, we use a larger class of test functions and we assume from the beginning that the solution is continuous in time.
@@
A map $m\in\mathcal{S}$ is called a weak solution to (3.1) if $m(0)=m_{0}$ and
@@
for every $t\in[0,T]$ and $\varphi\in\mathcal{D}_{T}$. [^1] Instead of the requirement that $m\in\mathcal{S}$, in [6] it was assumed that a weak solution $m:[0,T]\rightarrow\mathcal{P}_{1}(H)$ is such that $t\mapsto m(t)(I)$ is measurable for each $I\in\mathcal{B}(H)$, that is, in the terminology of [6], $m$ is a probability kernel.
@@
This is just an application of Dynkin’s formula for sufficiently regular functions: see, e.g., [21, Prop. 1.169], applied there with
$b_{0}(s,x,a_{1})=a_{1},\ a_{1}(s,\omega)=w(s,X^{w}_{s}(\omega)),\ a_{2}(\cdot)=0,\ \Lambda=B_{\|w\|_{\infty}}\subset H.\vspace{-.9cm}$
@@
The claim follows from [6, Th. 4.1] once we prove that a solution in our sense is also a solution in the sense of [6, Th. 4.1]. To do that we first notice that, in [6, Formula (1.5)], the definition of solution is given using the set test functions (called $\mathcal{E}_{A}(H)$ there) which is the linear span of the real parts of the functions of the form
@@
Given the above, we observe that such a set is clearly contained in the set $\mathcal{D}_{T}$ defined in (3). Finally, we point out that in [6] the Kolmogorov operator $L$ is defined differently (corresponding to our $\partial_{t}+L_{0}-\langle w,D\rangle$); still, equality (1.5) in [6] corresponds to our (3.4).
@@
By Proposition 2.5(i), our solution provided by Proposition 3.2 belongs to the class $\mathcal{S}_{4}$ and the uniqueness follows from Proposition 3.3.
@@ 4. The HJB Equation: Mild Solutions
In this section we deal with equation (HJB) for a given fixed $m\in\mathcal{S}$. We introduce the concept of a mild solution to (HJB) (see [21, Section 4.4.1.2]).
@@
    This part is contained in [21, Theorem 4.90]. [^2] More precisely, it is enough to take in that theorem, $m=0$, $\gamma_{G}(t)=t^{-\gamma}$, $G=\mathbf{Id}$. We warn that $G$ in this footnote refers to the operator used in [21, Theorem 4.90], which has nothing to do with our function $G$.
@@
    Using the smoothing properties of the Ornstein-Uhlenbeck semigroup (see e.g. [21, Theorem 4.37]), we have the representation of the gradients
@@
    Hence, fixing $t\in[0,T)$ and considering [21, Th. 4.37] and Assumption 2.6, we have, for some $C>0$,
@@
    By a suitable generalization of Gronwall’s Lemma (see [21, Prop. D.30]), we get
@@ 5. The MFG System: Existence of Solutions
    $\sup_{t\in[0,T],\,x\in H}\int_{0}^{t}\langle e^{(t-s)A}w(s,x),\mathbf{e}_{n}\rangle^{2}\mathrm{d}s\leq\alpha_{n}\cdot\sup_{s\in(0,T),\,x\in H}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\ \leq\ \alpha_{n}{\|w\|^{2}_{\infty}},$
    for every $n\in\mathbb{N}.$
@@
    $\int_{H}\sup_{t\in[0,T]}\langle e^{tA}x,\mathbf{e}_{n}\rangle^{2}m_{0}(\mathrm{d}x)=\int_{H}\sup_{t\in[0,T]}\langle x,e^{tA}\mathbf{e}_{n}\rangle^{2}m_{0}(\mathrm{d}x)$
@@
    $=\sup_{t\in[0,T],\,x\in H}\ \int_{0}^{t}e^{2(t-s)\lambda_{n}}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\mathrm{d}s$
    $\leq\sup_{s\in(0,T),\,x\in H}\langle w(s,x),\mathbf{e}_{n}\rangle^{2}\cdot\sup_{t\in[0,T]}\int_{0}^{t}e^{2(t-s)\lambda_{n}}\mathrm{d}s$
@@
Observe that, by Lemma A.1 and by [41, Remark 7.13 (ii)], the metric $\mathbf{d}_{1}$ metrizes on $\mathcal{R}_{m_{0}}$ the topology of weak convergence in $\mathcal{P}(H)$. In view of Proposition 5.2, we also define
@@
We will use the fact (see, e.g., [41, Rem. 7.5(i)] or [11, Cor. 5.4]) that for $\mu,\nu\in\mathcal{P}_{1}(H)$
@@
6. The MFG System: Uniqueness of Solutions
In this section, we establish the uniqueness of solutions for our MFG system (HJB)–(FP). It is well known, even in finite-dimensional state spaces (see [10, Rem. 1.4, p.32]), that uniqueness of a solution for this system cannot generally be guaranteed. A typical approach to ensure uniqueness is to impose certain monotonicity conditions on the data. One of the most common monotonicity conditions is the Lasry-Lions monotonicity condition (see, for example, [33, 34, 35] and [10, Section 1.3.2]). For other monotonicity conditions that ensure uniqueness in finite-dimensional MFG systems, readers can refer to [31, 37] and the references therein. Another typical feature of uniqueness results is the separability of the Hamiltonian function $\mathcal{H}$ (as assumed in Assumption 6.1(i) here). Well-posedness of MFG systems without separability has been studied in [37]. Here, we assume separability and use a variant of the Lasry-Lions monotonicity condition from [10, Theorem 1.4]. Our assumptions are as follows.
@@ 6. The MFG System: Uniqueness of Solutions
      The following is satisfied:
      $\int_{H}[F(x,\mu_{1})-F(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)>0\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{1}(H),\;\mu_{1}\neq\mu_{2}.$
@@
      $\int_{H}[G(x,\mu_{1})-G(x,\mu_{2})](\mu_{1}-\mu_{2})(\mathrm{d}x)=0\quad\Longrightarrow\quad G(\cdot,\mu_{1})=G(\cdot,\mu_{2}).$
@@
Uniqueness. The proof is an adaptation to our Hilbert space setting of a typical proof of uniqueness in finite dimensional spaces (see, e.g., [10, Theorem 1.4]).
@@
$\bar{v}(t,x)=$ $\ R_{T-t}[G(\cdot,m_{1}(T))-G(\cdot,m_{2}(T))](x)$
$-\int_{t}^{T}R_{s-t}[\mathcal{H}^{0}(\cdot,Du_{1}(s,\cdot))-\mathcal{H}^{0}(\cdot,Du_{2}(s,\cdot))](x)\mathrm{d}s$
$\ \ \ -\int_{t}^{T}R_{s-t}[F(\cdot,m_{1}(s))-F(\cdot,m_{2}(s))](x)\mathrm{d}s,$
@@
$-\int_{0}^{T}\left(\int_{H}\langle\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x)),\,D\varphi(s,x)\rangle\,\,m_{1}(s,\mathrm{d}x)-\int_{H}\langle\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x)),\,D\varphi(s,x)\rangle\,\,m_{2}(s,\mathrm{d}x)\right)\mathrm{d}s.$
@@
by Dominated Convergence Theorem. The convergence of the other terms follows directly by Dominated Convergence Theorem, since the functions involved are uniformly bounded. Therefore, passing to the limit in (6.3), we get
@@
    If case (a) of Assumption 6.1(iv) holds, we have $m_{1}=m_{2}$; then uniqueness of mild solutions of equation (HJB) gives $v_{1}=v_{2}$.
@@
    If case (b) of Assumption 6.1(iv) holds, we first use uniqueness of mild solutions of equation (HJB), which holds since
    $F(x,m_{1}(s))=F(x,m_{2}(s)),\ \ \ G(x,m_{1}(T))=G(x,m_{2}(T)),$
    to obtain $v_{1}=v_{2}$; then, we conclude that $m_{1}=m_{2}$ by uniqueness of weak solutions of (3.4) with $w(s,x):=\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x))$.
@@
    If case (c) of Assumption 6.1(iv) holds, we first obtain
    $\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x)),$
    $m_{1}(s)$ and $m_{2}(s)$ a.e. for $s\in[0,T]$. This implies $m_{1}=m_{2}$ by uniqueness of weak solutions of (3.4) with $w(s,x)=\mathcal{H}^{0}_{p}(x,Dv_{1}(s,x))$ (or equivalently with $w(s,x)=\mathcal{H}^{0}_{p}(x,Dv_{2}(s,x))$) and then conclude that $v_{1}=v_{2}$ using uniqueness of mild solutions of equation (HJB).